Hello, can you please help me with the following equations. I am totally lost and still cannot grasp this. Could you please show the steps to getting the answer as well as I will use this as my study references.Thank you so much.
1. 5x + 3 < -2
2. 6x – 5 greater then or equal to -17
3. -9x + 4 less then or equal to -14
4. 9( x – 3 ) greater then or equal to -9
5. 5(x – 2) greater then or equal to 3(x + 2)

Answer by rfer(16322)

you work these just like a regular equation.
1)
5x+3<-2
subtract 3
5x<-2-3
5x<-5
divide by 5
x<-1
2)
6x-5>=-17
add 5 to both sides
6x>=-17+5
6x>=-12
x>=-2
3)
-9x+4<=-14
subtract 4
-9x<=18
divide by -9
when dividing by a minus at this point the sign reverses
x>=-2
4)
9(x-3)>=-9
9x-27>=-9
add 27
divide by 9
9x>=+18
x>=2
5)
5(x-2)>=3(x+2)
5x-10>=3x+6
add 10
5x>=3x+6+10
subtract 3x
2x>=16
x>=8
I hope that helps.
Bob